\begin{thebibliography}{10}


\bibitem{milgram1967small}
S.~Milgram, ``Small world experiment,'' \emph{Psychology Today}, vol.~1, pp.
  61--67, 1967.

\bibitem{watts1998collective}
D.~J. Watts and S.~H. Strogatz, ``Collective dynamics of
  ‘small-world’networks,'' \emph{nature}, vol. 393, no. 6684, pp. 440--442,
  1998.

\bibitem{Barabasi509}
A.-L. Barab\'asi and R.~Albert, ``Emergence of scaling in random networks,''
  \emph{Science}, vol. 286, no. 5439, pp. 509--512, 1999.

\bibitem{PhysRevLett.85.5234}
R.~Albert and A.-L. Barab\'asi, ``Topology of evolving networks: Local events
  and universality,'' \emph{Phys. Rev. Lett.}, vol.~85, pp. 5234--5237, Dec
  2000.

\bibitem{Girvan7821}
M.~Girvan and M.~E.~J. Newman, ``Community structure in social and biological
  networks,'' \emph{Proceedings of the National Academy of Sciences}, vol.~99,
  no.~12, pp. 7821--7826, 2002.

\bibitem{Batty1418}
M.~Batty, ``A theory of city size,'' \emph{Science}, vol. 340, no. 6139, pp.
  1418--1419, 2013.

\bibitem{doi:10.1098/rsif.2015.0937}
H.~Youn, L.~M.~A. Bettencourt, J.~Lobo, D.~Strumsky, H.~Samaniego, and G.~B.
  West, ``Scaling and universality in urban economic diversification,''
  \emph{Journal of The Royal Society Interface}, vol.~13, no. 114, p. 20150937,
  2016.

\bibitem{Bettencourt7301}
L.~M.~A. Bettencourt, J.~Lobo, D.~Helbing, C.~K\"uhnert, and G.~B. West,
  ``Growth, innovation, scaling, and the pace of life in cities,''
  \emph{Proceedings of the National Academy of Sciences}, vol. 104, no.~17, pp.
  7301--7306, 2007.

\bibitem{Bettencourt1438}
L.~M.~A. Bettencourt, ``The origins of scaling in cities,'' \emph{Science},
  vol. 340, no. 6139, pp. 1438--1441, 2013.

\bibitem{doi:10.1098/rspa.2019.0742}
I.~Artico, I.~Smolyarenko, V.~Vinciotti, and E.~C. Wit, ``How rare are
  power-law networks really?'' \emph{Proceedings of the Royal Society A:
  Mathematical, Physical and Engineering Sciences}, vol. 476, no. 2241, p.
  20190742, 2020.

\bibitem{RevModPhys.74.47}
R.~Albert and A.-L. Barab\'asi, ``Statistical mechanics of complex networks,''
  \emph{Rev. Mod. Phys.}, vol.~74, pp. 47--97, Jan 2002.

\bibitem{Song2005}
Song, C.~Havlin, S.~Makse, and H.~A., ``Self-similarity of complex networks,''
  \emph{Nature}, vol. 433, p. 392, 2005.

\bibitem{West1677}
G.~B. West, J.~H. Brown, and B.~J. Enquist, ``The fourth dimension of life:
  Fractal geometry and allometric scaling of organisms,'' \emph{Science}, vol.
  284, no. 5420, pp. 1677--1679, 1999.

\bibitem{Tsallis1988}
C.~Tsallis, ``Possible generalization of boltzmann-gibbs statistics,''
  \emph{Journal of Statistical Physics}, vol.~52, 1988.

\bibitem{doi:10.1089/cmb.2006.13.810}
R.~Khanin and E.~Wit, ``How scale-free are biological networks,'' \emph{Journal
  of Computational Biology}, vol.~13, no.~3, pp. 810--818, 2006, pMID:
  16706727.

\bibitem{ABBASI2020110119}
M.~Abbasi, A.~Bollini, J.~Castillo, A.~Deppman, J.~Guidio, P.~Matuoka,
  A.~Meirelles, J.~Policarpo, A.~Ramos, S.~Simionatto, A.~Varona,
  E.~Andrade-II, H.~Panjeh, and L.~Trevisan, ``Fractal signatures of the
  covid-19 spread,'' \emph{Chaos, Solitons \& Fractals}, vol. 140, p. 110119,
  2020.

\bibitem{PhysRevLett.86.3200}
R.~Pastor-Satorras and A.~Vespignani, ``Epidemic spreading in scale-free
  networks,'' \emph{Phys. Rev. Lett.}, vol.~86, pp. 3200--3203, Apr 2001.

\bibitem{RevModPhys.87.925}
R.~Pastor-Satorras, C.~Castellano, P.~Van~Mieghem, and A.~Vespignani,
  ``Epidemic processes in complex networks,'' \emph{Rev. Mod. Phys.}, vol.~87,
  pp. 925--979, Aug 2015.

\bibitem{PhysRevLett.85.4626}
R.~Cohen, K.~Erez, D.~ben Avraham, and S.~Havlin, ``Resilience of the internet
  to random breakdowns,'' \emph{Phys. Rev. Lett.}, vol.~85, pp. 4626--4628, Nov
  2000.

\bibitem{PhysRevD.93.054001}
A.~Deppman, ``Thermodynamics with fractal structure, tsallis statistics, and
  hadrons,'' \emph{Phys. Rev. D}, vol.~93, p. 054001, Mar 2016.

\bibitem{https://doi.org/10.1046/j.1461-0248.2002.00364.x}
A.~Belgrano, A.~P. Allen, B.~J. Enquist, and J.~F. Gillooly, ``Allometric
  scaling of maximum population density: a common rule for marine phytoplankton
  and terrestrial plants,'' \emph{Ecology Letters}, vol.~5, no.~5, pp.
  611--613, 2002.

\bibitem{https://doi.org/10.1890/03-9000}
J.~H. Brown, J.~F. Gillooly, A.~P. Allen, V.~M. Savage, and G.~B. West,
  ``Toward a metabolic theory of ecology,'' \emph{Ecology}, vol.~85, no.~7, pp.
  1771--1789, 2004.

\bibitem{Deppman_2020}
A.~Deppman, E.~Meg\'ias, and D.~P. Menezes, ``Fractal structure of yang-mills
  fields,'' \emph{Physica Scripta}, vol.~95, no.~9, p. 094006, sep 2020.

\bibitem{PhysRevD.101.034019}
------, ``Fractals, nonextensive statistics, and qcd,'' \emph{Phys. Rev. D},
  vol. 101, p. 034019, Feb 2020.

\bibitem{newman2003structure}
M.~E. Newman, ``The structure and function of complex networks,'' \emph{SIAM
  review}, vol.~45, no.~2, pp. 167--256, 2003.

\bibitem{PhysRevE.87.062806}
G.~Bianconi, ``Statistical mechanics of multiplex networks: Entropy and
  overlap,'' \emph{Phys. Rev. E}, vol.~87, p. 062806, Jun 2013.

\bibitem{10.3389/fevo.2019.00242}
C.~P. Kempes, M.~A.~R. Koehl, and G.~B. West, ``The scales that limit: The
  physical boundaries of evolution,'' \emph{Frontiers in Ecology and
  Evolution}, vol.~7, p. 242, 2019.

\bibitem{LI2021105465}
B.~Li and X.~Li, ``A cross-diffusive evolution system arising from biological
  transport networks,'' \emph{Communications in Nonlinear Science and Numerical
  Simulation}, vol.~92, p. 105465, 2021.

\bibitem{FnetSocial}
H.~M. J, M.~B. T, W.~R. S, B.~Oskar, and B.~J. H, ``The complex structure of
  hunter–gatherer social networks,'' \emph{Proc. R. Soc. B.}, vol. 274, pp.
  2195--2203, 2007.

\bibitem{prima2018combining}
M.-C. Prima, T.~Duchesne, A.~Fortin, L.-P. Rivest, and D.~Fortin, ``Combining
  network theory and reaction--advection--diffusion modelling for predicting
  animal distribution in dynamic environments,'' \emph{Methods in Ecology and
  Evolution}, vol.~9, no.~5, pp. 1221--1231, 2018.

\bibitem{BiologicalNetwork-Nature}
X.~Liu, E.~Maiorino, A.~Halu, K.~Glass, R.~B. Prasad, J.~Loscalzo, J.~Gao, and
  A.~Sharma, ``Robustness and lethality in multilayer biological molecular
  networks transport networks,'' \emph{Nature Communications}, vol.~11, p.
  105465, 2021.

\bibitem{Keeling}
M.~J. Keeling and K.~T.~D. Eames, ``Networks and epidemic models,''
  \emph{Journal of the Royal Society, Interface}, vol.~2, pp. 295--307, 2005.

\bibitem{GONG2021125501}
H.~Gong, C.~Guo, and Y.~Liu, ``Measuring network rationality and simulating
  information diffusion based on network structure,'' \emph{Physica A:
  Statistical Mechanics and its Applications}, vol. 564, p. 125501, 2021.

\bibitem{2021PhyA..56125266F}
M.~Fu, J.~Feng, D.~Lande, O.~Dmytrenko, D.~Manko, and R.~Prakapovich, ``Dynamic
  model with super spreaders and lurker users for preferential information
  propagation analysis,'' \emph{Physica A Statistical Mechanics and its
  Applications}, vol. 561, p. 125266, jan 2021.

\bibitem{HAN2020125033}
W.~Han, Y.~Feng, X.~Qian, Q.~Yang, and C.~Huang, ``Clusters and the entropy in
  opinion dynamics on complex networks,'' \emph{Physica A: Statistical
  Mechanics and its Applications}, vol. 559, p. 125033, 2020.

\bibitem{Bettencourt19540}
L.~M.~A. Bettencourt and J.~Kaur, ``Evolution and structure of sustainability
  science,'' \emph{Proceedings of the National Academy of Sciences}, vol. 108,
  no.~49, pp. 19\,540--19\,545, 2011.

\bibitem{4118472}
R.~Olfati-Saber, J.~A. Fax, and R.~M. Murray, ``Consensus and cooperation in
  networked multi-agent systems,'' \emph{Proceedings of the IEEE}, vol.~95,
  no.~1, pp. 215--233, 2007.

% 
% 
% \bibitem{bib1}
% Conant GC, Wolfe KH.
% \newblock {{T}urning a hobby into a job: how duplicated genes find new
%   functions}.
% \newblock Nat Rev Genet. 2008 Dec;9(12):938--950.
% 
% \bibitem{bib2}
% Ohno S.
% \newblock Evolution by gene duplication.
% \newblock London: George Alien \& Unwin Ltd. Berlin, Heidelberg and New York:
%   Springer-Verlag.; 1970.
% 
% \bibitem{bib3}
% Magwire MM, Bayer F, Webster CL, Cao C, Jiggins FM.
% \newblock {{S}uccessive increases in the resistance of {D}rosophila to viral
%   infection through a transposon insertion followed by a {D}uplication}.
% \newblock PLoS Genet. 2011 Oct;7(10):e1002337.

\end{thebibliography}

\documentclass[10pt,letterpaper]{article}
\usepackage[top=0.85in,left=2.75in,footskip=0.75in]{geometry}

\usepackage{amsmath,amssymb}

\usepackage{changepage}

\usepackage[utf8x]{inputenc}

\usepackage{textcomp,marvosym}

\usepackage{cite}

\usepackage{nameref,hyperref}

\usepackage{microtype}
\DisableLigatures[f]{encoding = *, family = * }

\usepackage[table]{xcolor}

\usepackage{array}

\newcolumntype{+}{!{\vrule width 2pt}}

\newlength\savedwidth

\usepackage[aboveskip=1pt,labelfont=bf,labelsep=period,justification=raggedright,singlelinecheck=off]{caption}

\makeatletter
\renewcommand{\@biblabel}[1]{\quad#1.}
\makeatother

\usepackage{lastpage,fancyhdr,graphicx}
\usepackage{epstopdf}

\usepackage{cleveref}
\newtheorem{thm}{Theorem}[section]
\newtheorem{lem}{Lemma}[thm]
\newtheorem{prop}{Axiom}[section]
\newtheorem{cor}{Corollary}[thm]
\newtheorem{dfn}{Definition}[section]

\Crefname{thm}{Theorem}{Theorem}
\Crefname{lem}{Lemma}{Lemma}
\Crefname{prop}{Axiom}{Axiom}
\Crefname{cor}{Corollary}{Corollary}
\Crefname{dfn}{Definition}{Definition}

\begin{document}
\vspace*{0.2in}

\begin{flushleft}
{\Large
\textbf\newline{Flux of information in scale-free networks} 
}
\newline
\\
Airton Deppman\textsuperscript{1\Yinyang},
Evandro O. Andrade-II\textsuperscript{2\Yinyang},
\\
\bigskip
\textbf{1} Instituto de F\'isica, Universidade de S\~ao Paulo, S\~ao Paulo, SP, Brazil
\\
\textbf{2} Departamento de F\'isica, Universidade Estadual de Santa Cruz, Ilh\'eus, BA, Brazil
\\
\bigskip

%
%
\Yinyang AD developed the conceptual aspects and calculations. EOA-II worked on the calculations. Both authors contributed equally to the text.





* deppman@usp.br

\end{flushleft}
\section*{Abstract}
Scale-free networks constitute a fast-developing field that has already provided us with important tools to understand natural and social phenomena. From biological systems to environmental modifications, from quantum fields to high energy collisions, or from the number of contacts one person has, on average, to the flux of vehicles in the streets of urban centres, all these complex, non-linear problems are better understood under the light of the scale-free network's properties. A few mechanisms have been found to explain the emergence of scale invariance in complex networks, and here we discuss a mechanism based on the way information is locally spread among agents in a scale-free network. We show that the correct description of the information dynamics is given in terms of the q-exponential function, with the power-law behaviour arising in the asymptotic limit. This result shows that the best statistical approach to the information dynamics is given by Tsallis Statistics.  We discuss the main properties of the information spreading process in the network and analyse the role and behaviour of some of the parameters as the number of agents increases. The different mechanisms for optimization of the information spread are discussed.



\section*{Introduction}
A large number of problems that are common to modern societies can be addressed in the framework of complex networks. Accurate data and methods were made available by new technologies that are used worldwide, providing for the first time the adequate conditions to the development of scientific approaches to those problems. As a consequence, the last decades witnessed the fast evolution of our knowledge on the behaviour and properties of complex networks~\cite{milgram1967small, watts1998collective}.

Scale-free networks are of particular interest~\cite{Barabasi509,PhysRevLett.85.5234} since many aspects of physical, biological and sociological systems~\cite{Girvan7821,Batty1418,doi:10.1098/rsif.2015.0937, Bettencourt7301,Bettencourt1438} can be described to a good approximation by networks belonging to this class\cite{doi:10.1098/rspa.2019.0742}. Several mechanisms for the emergence of scaling symmetry in complex networks have been identified~\cite{RevModPhys.74.47}. Since scaling symmetry and a fine complex structure are the main features of fractals, this class of networks is also called a fractal network~\cite{Song2005}. The fractal dimension can be related to topological characteristics of the system~\cite{West1677}.

In this work, we describe and prove some of the most important characteristics of the flux of information in a fractal network. Information, here, is considered in a broad sense, and can refer to pieces of information locally transmitted, or to people or objects that move from one node to another in the network. We say the information is locally transmitted because we consider only those cases where the information is transmitted from one agent to a limited and small number of agents in the same network. These characteristics are present in several of our socioeconomic activities, and in natural systems. We discuss that the variables describing the different quantities in the system must appear in a scale-free form. In particular, we show that the time spent to share information among agents in the network is proportional to the squared-root of the number of agents. We argue that the scaling symmetry is broken at some point, and due to this symmetry break the information spreading follows a q-exponential function, and the statistical aspects of the network are hereby associated with the Tsallis Statistics~\cite{Tsallis1988}. 

The scale symmetry in complex networks is easily identified by observing the power-law behaviour of the investigated quantity distribution. Such distributions are found in biological systems~\cite{doi:10.1089/cmb.2006.13.810,Girvan7821}, as in the branching pattern of the circulatory system, in the metabolic rate of mammals~\cite{West1677}; in Epidemics~\cite{ABBASI2020110119,PhysRevLett.86.3200, RevModPhys.87.925}; Internet structure~\cite{PhysRevLett.85.4626}; in socioeconomic aspects of urban life, as in the number of patents or in the distribution of wages; in physical systems, as in high energy collisions, in thermodynamics systems~\cite{PhysRevD.93.054001}, Environmental Science~\cite{https://doi.org/10.1046/j.1461-0248.2002.00364.x, https://doi.org/10.1890/03-9000}  or in Yang-Mills fields theory~\cite{Deppman_2020, PhysRevD.101.034019}. Comprehensive reviews on the subject can be found in Refs~\cite{RevModPhys.74.47,newman2003structure}

Important advances in the statistical analysis of the scale-free networks have been made~\cite{RevModPhys.74.47,PhysRevE.87.062806}. On the other hand, the statistical aspects of fractal systems are associated with the Tsallis Statistics~\cite{Tsallis1988}, as it was shown in~\cite{PhysRevD.93.054001}. In the information dynamics studied here, we obtain the q-exponential behavior of the number of informed agents in the network. While the power-law behaviour appears in the asymptotic limit, the q-exponential function, which is typical of the Tsallis statistics, describes the full range of the distribution. The exponential behaviour is a limiting case when the number of edges per agent increases indefinitely. The work is organized as follows:
in~\Cref{scn:Scralefreenet} we describe the structure of the scale-free network, the scaling parameter and the break of the scaling symmetry; in~\Cref{scn:fluxinformation} we describe the information spreading dynamics over the scale-free network, obtain the q-exponential behaviour for the spreading and show that any variable must follow a power-law distribution. We obtain the differential equations that describe the information spreading dynamics and duscuss the different forms to increase the spreading efficiency. In~\Cref{scn:Discussion} we discuss the main results and the possibilities to apply this theory to different problems and in~\Cref{scn:Conclusions} we present our conclusions.

\section{Properties of the scale-free network}\label{scn:Scralefreenet}

Scale-free networks or, equivalently, fractal networks~\cite{Song2005} are represented by sets of nodes, connected among themselves. Each node in the network is itself a fractal network, similar to the initial one when its parameters are appropriately scaled. This hierarchical structure is a prominent characteristic of fractal networks, organizing the agents in an undetermined number of levels. It is always possible to find a scaling parameter which allows describing any node at any level of the fractal structure by the same mathematical expressions and in terms of scale-free variables.

These constraints on the network structure can be relaxed by the inclusion of probability distributions to determine some of the network features, but the distributions must be scale-free to preserve the scaling invariance. We will argue, in~\Cref{scn:Discussion}, that the results obtained here are very general and apply even in the case the constraints used here are relaxed.

In the following, we give definitions and derive some properties of a fractal network.



\begin{dfn}
 The fractal network is a set of $N$ nodes, also called agents, totally or partially connected to the other nodes in the set. This network of $N$ agents will be called main network. \label{dfn:2.1}
\end{dfn}

\begin{dfn}
 A scale-free, or fractal, network is a totally connected network. Each node is a fractal network similar to the main network, differing only by a scale parameter, forming an hierarchical structure. In this particular network, each agent is always comprised by $N$ agents in the next level. For simplicity we will refer to the fractal network also by fnet. \label{dfn:2.2}
\end{dfn}

\begin{dfn}
 If an agent $B$ is a component of an agent $A$, we say $B$ is an internal agent of $A$, and that $A$ is the parent agent of $B$.
\end{dfn}

\begin{prop}
 The agents of a fractal network are connected exclusively to the agents in the same network and to their parent agents. \label{prop:2.1}
\end{prop}

\begin{prop}
 All the properties of the agents are constrained to keep the similarity of the fractal network. \label{prop:2.2}
\end{prop}

\begin{thm}
 The fractal network has a natural scale, $\lambda$, associated with the total number of agents in the fractal network. \label{thm:2.1}
\end{thm}

\begin{lem}
 The fractal network presents a hierarchical structure. \label{lem:2.1}
\end{lem}
{\it Proof:} From~\Cref{dfn:2.2} we observe that the fnet is composed of $N$ agents interconnected, and each one is itself a fnet, therefore each agent has its own internal structure. Due to \Cref{prop:2.1}, the internal structure of each internal agent is not connected to any other agent outside its network. This property establishes a level structure in the fractal network ranked according to the number of internal structures one needs to consider until a specific node is reached.

\begin{dfn}
 We say an agent is at a level $l$ of the fractal structure if one needs to look into the internal structure of $l$ agents, starting from one of the agents in the main fractal network, in order to find that agent. Agents at the main fnet are at the level $l=1$. \label{dfn:2.3}
\end{dfn}

\begin{cor}
 The total number of agents at the level $l$ is $N^l$.
\end{cor}

{\it Proof:} Due to~\Cref{dfn:2.1} the fnet has $N$ agents at its first level. Due to~\Cref{dfn:2.2} each of those agents have $N$ internal agents, so at the second level, one has $N^2$. At each new level the number of agents is multiplied by $N$, so at the level $l$ the number of agents is $N^l$.

\begin{dfn}
 The quantities that characterize the agents are of two types: or they are constant and are called parameters, or they are variables. \label{dfn:2.4}
\end{dfn}

\begin{cor}
 Any variable, $\nu$ of the fnet must appear in the form $\nu/N^l$ and this ratio have to be scale invariant.
\end{cor}

{\it Proof:} Due to~\Cref{prop:2.2}, the variables must scale according to some quantity that characterizes the agent size. The number of internal agents is a natural quantity to be used as a scaling parameter.

\begin{dfn}
  The size of the agents of a fnet can be unequivocally set by defining a level $L$ at which the internal structure of the agents do not exist or can be disregarded in all relevant aspects of the fnet. The agents at this level are called individuals, and we refer to this level as the individual level.  \label{dfn:2.5}
\end{dfn}

\begin{dfn}
 The scaling parameter,  $\lambda_l$, that unequivocally determines the size of all agents in a fnet can be defined as  $\lambda_l=N^{L-l}$, where $L$ is the level of the individuals. \label{dfn:2.6}
\end{dfn}

\begin{lem}
 At the individual level the scaling symmetry is broken. \label{lem:1.2}
\end{lem}

{\it Proof:} This result follows immediately from~\Cref{dfn:2.5}, since the individual agents do not have an internal structure.

\begin{cor}
 The number of internal individuals in an agent at the level $l-1$ is given by $\sigma_{\lambda_l}=N\lambda_l$. \label{cor:2.1.3}
\end{cor}

{\it Proof:} It follows immediately from~\Cref{dfn:2.6}.

\begin{cor}
 Since the scaling symmetry is valid at any level, aside from the individual level, the whole fnet can be seen as a single agent. Attributing the level $l=0$ to the fnet allows one to represent the whole fnet as a single agent with $N^L$ internal individuals.
\end{cor}

{\it Proof:} It follows from~\Cref{dfn:2.6} and from~\Cref{cor:2.1.3}

This last result shows that there is a natural scale parameter, $\lambda$, that can be used to characterize the fnet, proving the~\Cref{thm:2.1}.

\section{Flux of information in a Fractal Network \label{scn:fluxinformation}}

In this section, we define what is meant by information and how it flows in the fractal network. The information is spread in the network from one initial agent that possess that piece of information, which is called an informed agent and transmits it or to the agents that are connected to it, or to its internal agents. When the piece of information reaches an uninformed agent it has a probability $\tau$ to received by that agent, and when it happens the uninformed agent becomes an informed one.

Information can be transmitted only by informed agents, and only uninformed agents can receive it. When an agent receives the piece of information, its first action is to transmit that piece of information to one of its internal agents. In the case of individual agents, it has a probability $\tau$ to change its state from uninformed to informed. When a fraction $\varphi$ of the internal agents are informed, the parent agent is considered informed.

Below we provide a formal description of the form of information spread in a fractal network and prove some of its characteristics. We also discuss some essential aspects of the scaling symmetry break, which will give rise to the q-exponential function and ultimately to the power-law behaviour.

\begin{dfn}
 Information is considered here in a broad sense, what includes pieces of information that can be shared among agents but also objects and people moving from one node to another. \label{dfn:3.1}
\end{dfn}

\begin{dfn}
 When an agent obtains the information, we say his states changes from uninformed to informed. The states of an agent are just informed or uninformed. \label{dfn:3.2}
\end{dfn}

\begin{prop}
 Agents can obtain the information only from informed agents with which they are connected, or from their parent agent. We say that, in this case, information is locally transmitted. \label{prop:3.1}
\end{prop}

\begin{prop}
  When an informed agent has a connection with an uninformed agent, we say that the information has reached the uninformed agent.   \label{prop:3.2}
\end{prop}

\begin{prop}
  When a piece of information reaches an uninformed agent, it is passed to its internal agents aleatory chosen.   \label{prop:3.3}
\end{prop}

\begin{prop}
  When a piece of information reaches an individual agent, it has a probability $\tau$ to accept that information. Upon acceptance, the individual becomes informed. The elapse of time to inform one individual is $\Delta t_o$.  \label{prop:3.4}
\end{prop}

\begin{prop}
  An agent that is not an individual becomes informed when a fraction $\tau$ of its internal agents are informed. \label{prop:3.5}
\end{prop}

\begin{prop}
  Uninformed agents cannot transmit information. Informed agents necessarily transmit information. \label{prop:3.6}
\end{prop}

\begin{thm}
 If $\lambda_L$ is the scale at the individual level and $\sigma$ is the total number of individuals in the fnet, the number of informed agents, $\nu$, after the start of information transmission is given in terms of a q-exponential function of $\sigma$ as
 \begin{equation}
  \nu(\sigma)=e_q(\tau \sigma)\,,
 \end{equation}
where
\begin{equation}
 e_q(\tau \sigma)=\left[ 1+(1-q) \frac{\tau \sigma}{\lambda_L} \right]^\frac{1}{1-q}\,,
\end{equation}
with
\begin{equation}
 1-q=1/N \,. \label{eqn:q}
\end{equation} \label{thm:3.1}
\end{thm}
The parameter $q$ is called q-index, and is completely determined by $N$.

\begin{lem}
 The number of informed agents after a period of information transmission at a level of the fractal network is given by
 \begin{equation}
 \nu(\tau)=(1+\tau)^{\alpha}\,, \label{eqn:pwr-lw}
\end{equation}
\end{lem}

{\it Proof:} According to~\Cref{prop:3.1}  and to~\Cref{prop:3.4}   the information is exchanged by $N$ agents where, initially, one of the agents is informed and all others are uninformed. An uninformed agent can get the information by different modes: it can get it directly from the first informed agent in the network, or it can get it from other agents that got the information from the initial agent, or even by more indirect ways. Mathematically, the number of informed agents after a period of information transmission is given by
\begin{equation}
 \nu(\tau)=1+\alpha \, \tau + (1/2)\,\alpha (\alpha-1) \tau ^2+\dots + C(\alpha,k)\tau ^k+ \dots + \tau ^{\alpha}\,, \label{eqn:PLexpansion}
\end{equation}
where $\alpha$ is the number of modes for the transmission of the information in the network. As one agent cannot transfer information to itself, we have that $\alpha=N-1$.

The combinatorial factor
\begin{equation}
 C(\alpha,k)=\frac{\alpha!}{(\alpha-k)! k!}
\end{equation}
arises because when the information is transferred to an uninformed agent, the order in which the informed agent obtained the information is not important. \Cref{eqn:PLexpansion} can be written as the power-law in~\Cref{eqn:pwr-lw}, proving this Lemma.

\begin{dfn}  
  We denominate $\sigma_{l,\lambda_{l'}}$, with $l'>l$, the number of internal agents with size, or scale, $\lambda_{l'}$ that an agent at the level $l$ has.
\end{dfn}

\begin{cor}  
 The number of informed agents can be expressed in terms of the scale parameter, $\lambda_l$.
\end{cor}

{\it Proof:} According to~\Cref{cor:2.1.3}, each agent in a level $l-1$ has exactly $\sigma_{l-1,\lambda_l}=N \lambda_l$ internal agents with size $\lambda_l$. Therefore we can write $\tau=\tau \sigma_{l-1,\lambda_l}/(N\lambda_l)$. It follows immediately from~\Cref{eqn:pwr-lw} that
\begin{equation}
 \nu(\tau)=\left(1+\tau \frac{\sigma_{l-1,\lambda_l}}{N\lambda_l}\right)^{N-1} \,. \label{eqn:quasiTsallis}
\end{equation}

\begin{cor}  
 The ratio $\sigma_{l,\lambda{l'}}/\lambda_{l'}$ is scale invariant.
\end{cor}

{\it Proof:} If you multiply $\lambda_{l'}$ by any positive, finite factor, due to~\Cref{cor:2.1.3} the number of agents at any level $l<l'$ is multiplied by the same factor, hence the ratio above remains invariant.

\begin{cor}  
 The symbols $\sigma_L$ and $\lambda_{L+1}$ are meaningless.
\end{cor}

{\it Proof:} According to~\Cref{dfn:2.6}, the level $L$ corresponds to that where agents are individuals, so they do not present an internal structure, hence there is no meaning in asking about its internal population. The individual size is the minimum size and determines the fundamental scale of the fnet, thereby there is no meaning in asking about scales below $\lambda_L$.

\begin{dfn}  
 When the scale is set to the individual size, that is, $\lambda_{l'}=\lambda_L$, we use the simplified notation $\sigma_l=\sigma_{l,\lambda_L}$. Accordingly, we denote by $\sigma$ the population of individuals in the fnet, that is, $\sigma=\sigma_0$.
\end{dfn}

\begin{cor}  
At the individual level we have $\lambda_L = 1$, the scale symmetry is broken and the q-exponential function is obtained.
\end{cor}

{\it Proof:} Adopting $\lambda_l=\lambda_L$, \Cref{eqn:quasiTsallis} becomes
\begin{equation}
  \nu(\tau)=\left(1+\tau \frac{\sigma}{N\, \lambda_L}\right)^{N-1}\,, \label{eqn:qexptau}
\end{equation}
where we used, for the sake of clarity, $\sigma= \sigma_{0,\lambda_L}$ as the total population of the network, recognized as a multiple the number of individuals in the network. Using the q-index defined by~\Cref{eqn:q} results that
\begin{equation}
  \nu(\sigma)=\left[1+ (1-q) \frac{\tau \sigma}{\lambda_L}\right]^{\frac{q}{1-q}}\,. \label{eqn:qexp}
\end{equation}
The expression above is not scale invariant, because now $\lambda_l=\lambda_L$ is fixed, and any variation of the fnet population, $\sigma$, results in a q-exponential behaviour.

This result proves the~\Cref{thm:3.1}. Notice that with the introduction of the individual level, indicated by the scale $\lambda_L$, the scale invariance disappears and we obtain according to the q-exponential function.

An additional comment is necessary at this point. Observe that the argument of the functions in~\Cref{eqn:qexptau} and~\Cref{eqn:qexp} are different. This results from the transition from a scale-free network to a fixed scale network. In the first case, the population increases according to the size of $\lambda_{l'}$ and the~\Cref{cor:2.1.3} is satisfied. In the second case, due to the symmetry break, the population increases while the scale is fixed. The number of close contacts agent is also fixed, as well as the parameter $q$. This means that the number of degrees of freedom for the information spread is independent of the population size. This is the main aspect of network for the emergence of non-extensivity, as will be discussed in~\Cref{scn:Discussion}

\subsection{The dynamics of the information spread}

In this section, we describe how to describe the dynamics of the information spread by including the time evolution of the number of informed agents.

\begin{dfn}
 We define the information spread time interval, $\Delta t_{\lambda}$, and the rate of transmission of information, $\kappa_{\lambda}$, such that $\tau=\kappa_{\lambda} \Delta t_{\lambda}$. \label{dfn:kappa}
\end{dfn}

\begin{thm}
  The time interval $\Delta t_{\lambda}$ depends on the population size,  $\sigma_{\lambda}$ as a power-law function, that is, $\Delta t_{\lambda}=\sigma^{\beta} \Delta t_{\lambda_L}$, and the transmission rate depends on the population size according to $\kappa_{\lambda}=\sigma_{\lambda}^{-\beta}\kappa_{\lambda_L}$.
 \label{thm:3.2}
\end{thm}

\begin{lem}
 The time interval of an agent, $\Delta t_{\lambda_l}$, is related to the internal agents time interval, $\Delta t_{\lambda_{l+1}}$ by  $\Delta t_{\lambda_l}=N^{\beta}\Delta t_{\lambda_{l+1}}$, with $0 \le \beta \le 1$. The rate of information transmission are related by $\kappa_{\lambda_l}=N^{-\beta}\kappa_{\lambda_{l+1}}$ \label{lem:3.1}
\end{lem}

{\it Proof:} Consider an agent at the scale $\lambda_l \ne \lambda_L$. According to~\Cref{dfn:2.1} this agent has $N$ internal agents. The elapsed time for the information transmission to an agent, $\Delta t_{\lambda_l}$, depends on the time its internal agents will demand to get the information, $\Delta t_{\lambda_{l+1}}$.

Due to \Cref{prop:2.2} all agents at the same hierarchic level corresponding to $\lambda_{l+1}$ have similar values for $\Delta t_{\lambda_{l+1}}$, so the maximum value for interval  for the parent agent is $\Delta t_{\lambda_{l}}=N \Delta t_{\lambda_{l+1}}$ in the case of the information is spread among the internal agents sequentially. The minimum value is $\Delta t_{\lambda_{l}}=\Delta t_{\lambda_{l+1}}$, in the case of simultaneous transmission of the information among the internal agents. In the gernal case we write $\Delta t_{\lambda_{l}}=N^{\beta}\Delta t_{\lambda_{l+1}}$, with $0 \le \beta \le 1$. The equalities correspond to the two special cases mentioned above. As $\tau$ is a parameter, according to \Cref{dfn:2.5} we must have $\kappa_{\lambda}=N^{-\beta}\kappa_{\lambda'}$. \label{lem:3.2.1}

\begin{cor}
 The parameter $\beta$ is independent of the agent level.
\end{cor}

{\it Proof:} It follows from the~\Cref{prop:2.2}.

\begin{cor}
 The variables $\Delta t_{\lambda_1}$ of an agent at level $l_1$ is related to the variable $\Delta t_{\lambda_2}$ at the level $l_2>l_1$ by $\Delta t_{\lambda_1}=N^{\beta(l_2-l_1)}\Delta t_{\lambda_2}$, and the variable $\kappa_{\lambda_1}$ is related to $\kappa_{\lambda_2}$ by $\kappa_{\lambda_1}=N^{-\beta(l_2-l_1)}\kappa_{\lambda_2}$. \label{lem:3.2.2}
\end{cor}

{\it Proof:} Applying recursively the result of~\Cref{lem:3.1} we get
\begin{equation}
 \Delta t_{\lambda_1}=\left(\prod_{i=l_1}^{l_2} N_i^{\beta} \right) \Delta t_{\lambda_2}=N^{\beta(l_2-l_1)}\Delta t_{\lambda_2}\,. \label{eqn:Deltat-arb}
\end{equation}
The result for $\kappa_{\lambda_1}$ can be obtained in the same way or by considering that $\tau$ is constant, as done in~\Cref{lem:3.2.1}.

Considering the result of \Cref{lem:3.2.2} for an agent at level $l_1=l$ and another agent at level $l_2=L$, we get
\begin{equation}
 \begin{cases}
  & \Delta t_{\lambda_l}=N^{\beta(L-l)}\Delta t_{\lambda_L} \\
  & \kappa_{\lambda_l}=N^{-\beta(L-l)}\kappa_{\lambda_L}
 \end{cases} \,.
\end{equation}
Setting $l=0$ we have $N^{L}=\sigma$, that is the number of individuals in the fnet, these results prove~\Cref{thm:3.2}.

\begin{thm}
 For a random fnet $\beta=0.5$. \label{thm:3.3}
\end{thm}

{\it Proof:} The time interval, $\Delta t$, for the information spread for an agent with a number $\sigma$ of internal agents if $\delta t$, is formed by the superposition of the intervals $\delta t$ for the information transmission to each of the internal agents.

If the interval of transmission for $n-1$ of the agents is $\Delta t_{n-1}$, the inclusion of an additional agent might increase the total time spent for information transmission only if the transmission in the $n$th agent starts at the instant $t$ such that $\Delta t- \delta t<t<\Delta t$.  This condition is satisfied with a probability $\delta t/\Delta t_n$, and when it happens the increase in the total interval of time is $\delta t/2$, on average, otherwise, the increase is null. Therefore we have, for $\sigma$ sufficiently large,
\begin{equation}
 \sigma \frac{d \Delta t}{d n}=\frac{1}{2}\sigma \frac{\delta t}{\Delta t} \delta t \,.
\end{equation}
If $\eta=n/\sigma$, we have $0\le \eta \le 1$, and the equation above becomes
\begin{equation}
 \frac{d \Delta t}{d \eta}=\frac{1}{2}\sigma \frac{\delta t}{\Delta t} \delta t\,.
\end{equation}
Integrating from $\eta=0$ to $\eta=1$ we have
\begin{equation}
 \Delta t=\sigma^{0.5} \delta t \,,
\end{equation}
what proves that $\beta=0.5$.

\subsection{Differential Equations for the information spread}

In this section we derive the differential equations governing the dynamics of the information spreading. In what follows we assume that, at any time, the number of agents being informed is much smaller than the total population, therefore the variation of the uninformed population during the elapse of time necessary to the newly informed agents change their states from uninformed to informed is negligible. This can be expressed mathematically by assuming that $\dot u t \ll u$ at any time.

\begin{dfn}
 The number of uninformed individuals in a population of individuals, $u(t)$, varies along time as more individuals receive the information and become informed agents. The uninformed population at any time is given by $u(t)=\sigma-\nu(t)$, where $\sigma$ is the population in the fnet, which is considered constant.. \label{dfn:3.5}
\end{dfn}

\begin{thm}  
 Given a small time interval $\delta t$, it is always possible to find an agent for which the elapsed time to spread the information is $dt < \delta t$. 
\end{thm}

{\it Proof:} Consider an arbitrary agent at a level $l_1$ whose spreading time is $\Delta t_{\lambda_1}$. If $\Delta t_{\lambda_1} < \delta t$, the condition is satisfied and the theorem is proved. If $\Delta t_{\lambda_1}> \delta t$, using~\Cref{eqn:Deltat-arb}, one can find a level $l_2>l_1$ at which the agents have a spreading time interval $\Delta t_{\lambda_2}$ such that
\begin{equation}
    \Delta t_{\lambda_2}= N^{-\beta(l_2-l_1)}\Delta t_{\lambda_1}<\delta t \,. \label{eqn:smooth}
\end{equation}

\begin{dfn}  
 We call smooth information spreading dynamics the process for which the individual spreading time is sufficiently small, so that for any {\it reasonably small} time interval $\delta t$, the elapsed time for an individual receive a piece of information, once the individual is reached by the spreading dynamics, is $\Delta t_{\lambda_L}<\delta t$.
\end{dfn}

In what follows we assume the spreading dynamics is smooth.

\begin{thm}  
 If at time $t$, measured in an appropriate scale for the dynamics of information at the individual level, a fnet has $u(t)$ uninformed individuals, the rate of increase in the number of informed individuals, $i(t)$, is represented by~\Cref{eqn:quasiTsallis}, which we write as
 \begin{equation}
  \frac{d i}{dt}=\kappa \frac{u(t)}{\lambda_L}\left[1+ (1-q) \frac{\tau u}{\lambda_L}\right]^{\frac{q}{1-q}}\,. \label{eqn:i-rate}
\end{equation}
\end{thm}

{\it Proof:} When a piece of information reaches an agent at the level $L-1$, it is passed to its $N$ internal individuals. In a population, $u$, of uninformed individuals, the number of agents at this level is
\begin{equation}
    M=\frac{u}{\lambda_{L}}\,,
\end{equation}
because of~\Cref{cor:2.1.3}. The number of those groups that receive the piece of information in the interval $dt$ is
\begin{equation}
    dM=\kappa M dt \,.
\end{equation}
For each group reached by the information, the number of individuals turning to the state informed is given by~\Cref{eqn:pwr-lw}, thus the number of individuals changing their state fron uninformed to informed is given by
\begin{equation}
    di=\kappa M dt \left(1+\tau\right)^{N-1} \,.
\end{equation}
Using~\Cref{eqn:q} and ~\Cref{dfn:3.5} we obtain~\Cref{eqn:i-rate}, proving the theorem.

\begin{cor}  
 The number of informed individuals in the network as a function of time is
 \begin{equation}
  i(t)=\left[1+(1-q) \frac{\kappa u(t) t}{\lambda_L}\right]^{\frac{1}{1-q}} \,, \,. \label{eqn:ifunc}
 \end{equation} \label{cor:3.5.1}
\end{cor}

{\it Proof:} Deriving~\Cref{eqn:ifunc} and using the assumption $\ddot u t \ll u$, we obtain the differential equation given in~\Cref{eqn:i-rate}, proving the theorem.

\begin{cor}  
 If a network is formed by $i_o$ fnets with independent spreading dynamics, the number of informed individuals is
 \begin{equation}
  i(t)=i_o\left[1+(1-q) \frac{\kappa u(t) t}{\lambda_L}\right]^{\frac{1}{1-q}}  \,,. \label{eqn:multi-ifunc} \,.
 \end{equation}
\end{cor}

{\it Proof:} It follows directly form~\Cref{cor:3.5.1}.

\begin{thm}
 The spread of information in a scale-free network can be described by the two coupled differential equations below:
 \begin{equation}
 \begin{cases}
  & \dfrac{d i(t_{\lambda_L})}{dt_{\lambda_L}}=\frac{i_o^{1-q}\kappa}{\lambda_L}i^q(t_{\lambda_L})\,u(t_{\lambda_L}) \\
  & \\
  & \dfrac{d u(t_{\lambda_L})}{dt_{\lambda_L}}=-\frac{i_o^{1-q}\kappa}{\lambda_L}i^q(t_{\lambda_L})\,u(t_{\lambda_L}) \\
 \end{cases}
\end{equation} \label{thm:3.5}
where $t_o$ is the instant when the spread of information starts.
\end{thm}

{\it Proof:} By differentiating~\Cref{eqn:multi-ifunc} we obtain the first equation above. Considering that the uninformed population is determined according to~\Cref{dfn:3.5}, the second equation is obtained.

\begin{thm}
 The solution to the coupled equations are
\begin{equation}
\begin{cases}
    & i(t_{\lambda_L})=i_o \,\left[1+(1-q)\frac{\kappa u(t)(t_{\lambda_L}-t_o))}{\lambda_L}(t_{\lambda_L}-t_o)\right]^{1/(1-q)}\, \\
    & u(t_{\lambda_L})=u(t_o)- i(t_{\lambda_L}-t_o) 
\end{cases} 
\end{equation} \label{pop-equations}
\end{thm}

\begin{thm}
 An approximate analytical solution for $u(t)$ can be  obtained, resulting in
\begin{equation}
  u(t)(\theta_{\lambda_L})=u(t_o)\left[1+ (1-q')\frac{\theta_{\lambda_L}-\theta_o}{\lambda'}  \right]^{-q'/(1-q')} \,, \label{eqn:u}
\end{equation}
with
\begin{equation}
\begin{cases}
    & q'=1-q \\
    & \theta_{\lambda_L}=\left(\frac{\kappa t_{\lambda_L} }{\lambda_L}\right)^{1/q'} \\
    & \lambda'=\frac{u(t_o)^{-(1-q')/q'}}{i_o} \,.
\end{cases} \label{eqn:variables}
\end{equation}
\end{thm}

{\it Proof:}

An approximate solution can be easily obtained by noticing that, in most cases of interest, we can approximate the equation for $i(t)$ by 
\begin{equation}
 \frac{d i(t)}{dt_{\lambda_L}}=-i_o\left[(1-q)^q\frac{\kappa (t_{\lambda_L}-t_o)}{\lambda_L}\right]^{1/(1-q)} i^{1/(1-q)}\, 
\end{equation}
that is a separable equation resulting in
\begin{equation}
 \frac{di(t_{\lambda_L})}{i^{1/(1-q)}}=-i_o\left[(1-q)^q\frac{\kappa t_{\lambda_L}}{\lambda_L}\right]^{1/(1-q)} dt_{\lambda_L}\,.
\end{equation}
Integrating both sides we get
\begin{equation}
 \int_{i_o}^{i}\frac{di'(t_{\lambda_L})}{i'^{1/(1-q)}}=-i_o\left[(1-q)^q\frac{\kappa }{\lambda_L}\right]^{1/(1-q)} \int_{t_o}^{t_{\lambda_L}} t'^{1/(1-q)} dt'\,.
\end{equation}
This equation results in
\begin{equation} \small
 i(t_{\lambda_L})^{-q/(1-q)}=i_o^{-q/(1-q)}+i_oq\left[\left(\frac{\kappa t_{\lambda_L} }{\lambda_L}\right)^{1/(1-q)}-\left(\frac{\kappa t_o }{\lambda_L}\right)^{1/(1-q)} \right]\,.
\end{equation}

Observe that with the definitions given in the~\Cref{eqn:variables} the equation above  can be conveniently written as~\Cref{eqn:u}, proving the theorem.

\subsection{Strategies for information diffusion}

One of the most important results of the investigation of flux of information through networks is the possibility to understand the optimization of the information spread dynamics, what  is important both for increasing the efficiency of communication and for formulating the best methods to avoid the information spread.

The main characteristic of the dynamics of information spread in the fractal network studied here is the local transmission of information by a small number of agents with close contact. The question that arises is the following: what is the best way to increase the efficiency of the information spread?

Two mechanisms could be devised to increase the efficiency: improve the probability of transmission, described by the parameters $\tau$ or by $\kappa$, or increasing the number of contacts between agents, given by $q$. We will see that the second option, when available, is the most effective.

\begin{thm}
 The increase of the rate of information transmission by increasing the efficiency of transmission is given by
 \begin{equation}
  i_q(\tau+\delta \tau)=\left[1+\frac{u/\lambda_L}{1+(1-q)\tau u/\lambda_L}\delta \tau \right]i_q(\tau)
 \end{equation}
\end{thm}

 {\it Proof:} Using~\Cref{dfn:kappa} in~\Cref{eqn:ifunc} and deriving with respect to $\tau$ we have
 \begin{equation}
  \frac{di_q}{d\tau}(\tau)=i_q^q(\tau) \,,
 \end{equation}
therefore the infinitesimal variation in the number of informed agents when the transmission probability varies from $\tau$ to $\tau+\delta \tau$ is
\begin{equation}
 \delta i_q(\tau)=i_q(\tau)\frac{u/\lambda_L}{1+(1-q)\tau u/\lambda_L}\delta \tau
\end{equation}
Hence, when the transformation $\tau \rightarrow \tau'=\tau+\delta \tau$ is performed, the number of informed agents transformation is
\begin{equation}
 i_q(\tau) \rightarrow i_q(\tau+\delta tu)=\left[1+\frac{u/\lambda_L}{1+(1-q)\tau u/\lambda_L}\delta \tau\right]i_q(\tau)\,. \label{eqn:tautrasnlationexact}
\end{equation}

\begin{cor}
 In the limit $\tau u/\lambda_L \gg 1/(1-q)$, the transformation of $\tau$ leads to a logarithmic increase in the number of informed agents.
\end{cor}

{\it Proof:} In this limit we have
\begin{equation}
 \frac{u/\lambda_L}{1+(1-q)\tau u/\lambda_L}\delta \tau \sim \frac{\delta \tau}{1+(1-q)\tau }=\frac{1}{1-q}\delta \log \tau \,.
\end{equation}
Substituting the result above in~\Cref{eqn:tautrasnlationexact} we obtain the logarithmic increase of the number of informed agents, i.e.,
\begin{equation}
 i_q(\tau) \rightarrow i_q(\tau+\delta \tau)=\left(1+\frac{1}{1-q}\delta \log \tau \right)i_q(\tau) \,.
\end{equation}

\begin{thm}
 The increase of the rate of information transmission by increasing the number of links per agent is
 \begin{equation}
  i_{q-\delta q}(\tau)=\left(1+\frac{q}{(1-q)^3}\delta q\right)i_{q-\delta q}(\tau)
 \end{equation} \label{cor:3.10}
\end{thm}
 {\it Proof:} Using~\Cref{dfn:kappa} in~\Cref{eqn:ifunc}, and deriving with respect to $N$ we have
 \begin{equation}
  \frac{d i_N(\tau)}{d N}=(1+N) \delta N \, i_N(\tau)\,.
 \end{equation}
From~\Cref{eqn:q} it follows that
\begin{equation}
 \frac{d i_q(\tau)}{d q}=-\frac{q}{(1-q)^3}i_q(\tau)
\end{equation}

From the results above we obtain that, under the transformation $q \rightarrow q-\delta q$, the number of informed agents transforms as
\begin{equation}
 i_{q}(\tau)\rightarrow i_{q-\delta q}(\tau)=\left[ 1+\frac{q}{(1-q)^3} \delta q \right]i_q(\tau)\,.
\end{equation}

\begin{cor}
The increase in the number of informed agents increases with $N^2$.
\end{cor}

{\it Proof:} Using~\Cref{thm:3.1} for the relation between $N$ and $q$, and the~\Cref{cor:3.10} we obtain
\begin{equation}
 i_{N}(\tau)\rightarrow i_{N+\delta N}(\tau)=\left[ 1+N \delta N \right]i_q(\tau)\,.
\end{equation}

\section{Discussion of the results} \label{scn:Discussion}

The characteristics of the network presented in~\Cref{scn:Scralefreenet} lead to the formation of a hierarchical scale-free structure typical of scale-free, or fractal, networks. The scaling parameter, $\lambda$, is given in terms of the number of internal agents. At some points the agents are considered as individuals with no internal structure, and at this point the scaling symmetry is broken.

The locally transmitted information and its spreading dynamics is defined in~\Cref{scn:fluxinformation}. The information is always shared among a number of connected agents in the same network or with the parent agent. This number is limited and constant throughout the network. This characteristic of the information spread dynamics and the broken scaling symmetry of the network structure give rise to a q-exponential function that describes the flux of information in the network. The power-law behaviour is obtained asymptotically, as the number of agents in the network increases.

The fact that the number of informed agents is described in terms of a q-exponential function, with the power-law behaviour obtained in the asymptotic limit, is an interesting result and deserves some additional comments. The q-exponential function results from the fact that the number of degrees of freedom of the spreading dynamics~\cite{PhysRevLett.85.5234} is uncorrelated to the number of agents in the network. This aspect of the fractal structure allows the number of individuals increase without any change in the number of modes by which an arbitrary agent can exchange information with another in the fnet. It is easy to understand, from the results obtained here, why fnets can describe so many aspects of natural and social systems: in many cases the information is transmitted locally among a small group of agents, and this number will be the same, no matter how many individuals one are in that population of the network.

The Theorems proved in~\Cref{scn:fluxinformation} show that any variable describing some quantity related to the information spread must appear in a scale-free form. The time interval for the spread of information in the network, for instance, increases with the squared-root of the number of individuals in the network. This result is in agreement with the Literature~\cite{RevModPhys.74.47,newman2003structure}, where the close connections between power-law distributions and scale-free networks is observed. In the present case, we show that any fractal network will depend only on power-law variables.

The rate of information spread given by the q-exponential function shows that the spread dynamics results in a slower transmission of information than one would expect in an exponential spread. But as $q \rightarrow 1$ the number of links among agents increases and the exponential behaviour is recovered. This corresponds to broadcast information, with chaotic transmission of information to all the individuals in the network. We verified that the most efficient way to increase the number of informed agents is not by increasing the transmission probability, but by increasing the number of connections among agents.

This result is interesting in many aspects, but here we would like to emphasize one of them. A virus undergoes random mutation, and the dominant strain will be more likely the one that can be transmitted more effectively. The way mutagenesis of virus can lead to a more effective spread is not by increasing its probability of transmission, which we associate with the parameter $\tau$, but by increasing the number of susceptible individuals in contact with the infected individuals, which we associate with $N$. Those strains that succeed to increase $N$ will be more effective in transmission, and therefore will be dominant. Thus, viruses will increase the multiplication factor more efficiently if they succeed to provide a longer  transmission time before the symptoms of its associated disease become evident.

As mentioned in the introduction, the definitions given in~\Cref{scn:Scralefreenet} and in~\Cref{scn:fluxinformation} can be relaxed in many ways. For instance, the number of internal agents can be set as variable, but must follow the same distribution whatever is the fractal level of the parent agent, and the corresponding variable must be scale-free, i.e., it must appear as fractions of the scaling parameter.
The same reasoning applies to the number of edges linking the agents. The information spread can follow an arbitrary distribution instead of being completely random, as far as the distribution is scale-free. Even the number of modes, or degrees of freedom, by which  the agents can obtain information from the others in the same network do not need to be constant. If these modifications are introduced in the scale-free network presented here, as far as the scale symmetry is preserved, our conclusions should hold.. Even if different numbers of edges among the nodes are used throughout the network, a multifractal network may be obtained. In all these cases, however, the general results obtained here will remain valid.

\section{Conclusion} \label{scn:Conclusions}

In this work we studied the spreading dynamics of information locally transmitted  through nodes, or agents, in self-similar, or fractal, network. The fractal network is defined by its fine internal structure that is scale-free. The self-similarity is a consequence of the the scaling property of the network and of its fine internal structure. However, at some point the scale symmetry is broken, and as a result the flux of information follows a q-exponential function, typical of the Tsallis' statistics. The pure power-law behavior results in the asymptotic limit, when the number of informed agents in the network is large. The exponential behavior, on the other hand, is obtained in the asymptotic limit of the number of connection of each agent increasing indefinitely.

The locally transmitted information, which goes from one agent to its neighbours and involves a limited number of nodes, independent of the total number of agents in the network, was studied. Its spreading dynamics reveal that the number of informed agents increases according to a q-exponential function. From the statistical point of view, this result indicates that the Tsallis Statistics is the correct framework to investigate the scale-free network. The constraint on the number of contact of each agent implies in an increasing number of levels in  the network as the population increases. This is contrary to the small-word hypothesis~\cite{milgram1967small,watts1998collective}, where the number of levels in the network is fixed and the number of contacts increases.

The time interval for the transmission increases as the squared-root of the number of individuals in the network. The scaling properties, establishes a power-law condition to the probability of transmission of a piece of information by agent in the network. Our results can be easily tested in real or simulated data by checking the characteristics of the distributions.

Differential equations describing the information spread are derived. We discussed the different strategies one can take to increase the information spread efficiency. These strategies can be formulated by increasing the transmission efficiency or by the number of connected agents. We show that the most effective strategy is the second one.

The results obtained in the present work have an impact on the formulation of the best strategies for the information spread. In practice, it may have implications on Environment Sciences~\cite{10.3389/fevo.2019.00242}, since modifications in the local environment may evolve by locally transmitted effects to larger and distant areas~\cite{LI2021105465,FnetSocial,prima2018combining}; Epidemiology, since viruses may evolve, by mutagenesis, to 
variants that optimize its transmission, a track that would prefer increasing the number of degrees of freedom by extending the period of virus transmission rather than increasing its transmission rate~\cite{BiologicalNetwork-Nature,Keeling}; Sociology~\cite{doi:10.1098/rsif.2015.0937}, since communication among individuals in the society can be made more or less effective by controlling the mechanisms of spreading, what may have an impact in policies and strategies to, e.g., combat fake-news and other irrational behaviours in social media~\cite{GONG2021125501, 2021PhyA..56125266F,HAN2020125033,Bettencourt19540}.  In Computer Sciences~\cite{PhysRevLett.85.4626}, Biology~\cite{doi:10.1089/cmb.2006.13.810}, Physics~\cite{PhysRevE.87.062806}, Economics~\cite{Bettencourt7301}, and Machine Learning~\cite{4118472}, to name just a few. 

The model of fractal network studied here can be modified in some aspects without changing the conclusions.

\section*{Acknowledgments}
AD would like to thank the support from the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under grant 304244/2018-0, and from by Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) under Grant No. 2016/17612-7. 


%
%
%

\end{document}